### v1

# On the Fragility of Active Learners

Abstract
Active learning (AL) techniques aim to maximally utilize a labeling budget by iteratively selecting instances that are most likely to improve prediction accuracy. However, their benefit compared to random sampling has not been consistent across various setups, e.g., different datasets, classifiers. In this empirical study, we examine how a combination of different factors might obscure any gains from an AL technique.
Focusing on text classification, we rigorously evaluate AL techniques over around $1000$ experiments that vary wrt the dataset, batch size, text representation and the classifier. We show that AL is only effective in a narrow set of circumstances. We also address the problem of using metrics that are better aligned with real world expectations.
The impact of this study is in its insights for a practitioner: (a) the choice of text representation and classifier is as important as that of an AL technique, (b) choice of the right metric is critical in assessment of the latter, and, finally, (c) reported AL results must be holistically interpreted, accounting for variables other than just the query strategy.
Keywords:
First keyword Second keyword Another keyword.
1 Introduction
Within a supervised learning setup, Active Learning (AL) techniques use a Query Strategy (QS) to identify an unlabeled set of instances which is optimal in the following sense: if labelled and added to the training data, they lead to the greatest improvement in model accuracy, relative to any other same-sized set. In cases where labelling is expensive, using AL is cost-efficient compared to random sampling, because the model reaches greater accuracy with a smaller number of labelled instances. AL techniques can be both domain-agnostic, e.g., uncertainty-based sampling [16], or domain specific, e.g., Contrastive Active Learning [21] for Natural Language Processing (NLP).
Our focus here is AL for text classification. Today, a practitioner has easy access to a variety of powerful classifiers via packages such as scikit-learn [25] and Hugging Face [34], and text representations, such as Universal Sentence Encoding (USE) [3] and MPNet [31]. Within that context, we ask these questions: is the choice of AL algorithm still as important? Do AL techniques exhibit strong performance over random sampling with any combination of text representation and a classifier?
In this empirical work, we answer this question in the negative: AL techniques can be fragile, i.e., they are heavily influenced by other variables. This study is valuable for the following reasons:
  1. 1.
    In practice, where data labeling is costly, should one just spend time seeking a good AL algorithm? Or should they cast their net of design choices wider to take into account the classifier and representation as well?
  2. 2.
    An important part of assessing a technique is how we measure, i.e., metrics used. We discuss this aspect too, e.g., rank-based metrics such as the Wilcoxon test or Mean Rank ignore how large is the accuracy gap wrt random sampling.
    The practical implication of this is that a small gap might not justify the cost of an AL workflow, given that random sampling is trivial to use in most programming languages.
  3. 3.
    Finally, this cautions us as consumers of research that an AL technique is not just the QS; reported results must be carefully understood to be situated within, and affected by, other significant experiment variables such as the classifier, representation, batch size, etc.
To our knowledge, this is the first such comprehensive study of AL algorithms for the task of text classification. The rest of the paper is organized as follows:
2 Previous Work
The field of AL has seen a steady stream of novel contributions over the years, e.g., [16, 12, 29, 7]. Being a hard problem - identifying important instances having only seen a few examples from the instance-label distribution - it also seems to suffer from real-world challenges, e.g., wrt variations in dataset, batch sizes, classifier used, which has been pointed out in the literature [1, 35, 18, 20].
In the domain of NLP, the problem is compounded by the fact that text representations may vary widely. This is pertinent since a lot of rich representations have been proposed in the past few years, e.g., word2vec [22], Universal Sentence Encodings (USE) [3], Backpacks [14], BERT [10], MiniLM [33], MPNet [31]. We are not aware of any work that studies AL’s effectiveness in context of these relatively modern representations. Hence, this work constitutes a useful addition to the line of work that draws attention to challenges of AL deployment.
3 Batch Active Learning - Overview
In this work, we study AL techniques for the task of text classification, in the batch AL setting. To ensure no ambiguity, we provide pseudo-code for this AL setting in Algorithm 1.
Input: Unlabeled data $X_{U}$, test data $(X_{test},Y_{test})$, query strategy $\mathcal{Q}$, seed set selection strategy $\mathcal{A}$, search space $\Theta$ for model $\mathcal{M}$, seed size $s$, batch size $b$, number of iterations $T$, metric $\mathcal{V}$
Result: Scores on test set at various iterations $\{(\mathcal{V}_{0},0),(\mathcal{V}_{1},1),...,(\mathcal{V}_{T},T)\}$
$result\leftarrow\{\}$ // to be returned
1 $X_{L,0},X_{U,0}\leftarrow\mathcal{A}(X_{U},s)$
2 $(X_{L,0},Y_{L,0})\leftarrow\text{obtain labels for }X_{L,0}$
$M_{0}\leftarrow\operatorname*{arg\,max}_{\theta\in\Theta}M_{\theta}((X_{L,0},Y_{L,0}))$ // both model selection and calibration are performed
3 $\mathcal{V}_{0}\leftarrow\mathcal{V}(M_{0}(X_{test}),Y_{test})$
4 $result\leftarrow result\cup\{(\mathcal{V}_{0},0)\}$
5 for $t\leftarrow 1$ to $T$ do
6       $X_{L,t}^{new},X_{U,t}\leftarrow\mathcal{Q}(M_{t-1},X_{U,t-1},(X_{L,t-1},Y_{L,t-1}),b)$
7       $(X_{L,t}^{new},Y_{L,t}^{new})\leftarrow\text{obtain labels for }X_{L,t}^{new}$
8       $(X_{L,t},Y_{L,t})\leftarrow\text{add }(X_{L,t}^{new},Y_{L,t}^{new})\text{ to previous labeled pool }(X_{L,t-1},y_{L,t-1})$
       $M_{t}\leftarrow\operatorname*{arg\,max}_{\theta\in\Theta}M_{\theta}((X_{L,t},Y_{L,t}))$ // both model selection and calibration are performed
9       $\mathcal{V}_{t}\leftarrow\mathcal{V}(M_{t}(X_{test}),Y_{test})$
10       $result\leftarrow result\cup\{(\mathcal{V}_{t},t)\}$
11
12 end for
return $result$
Algorithm 1 Batch Active Learning.
At a high-level, at every AL iteration $1\leq t\leq T$, we use a query strategy $\mathcal{Q}$ to select a $b$-sized batch of instances from the unlabeled pool of data (line 8). We obtain labels for this set (line 9) and add it to the existing pool of labeled data (line 10). We then train a model $M_{t}$ over this data (line 11). We emphasize that:
  1. 1.
    The model $M_{t}$ is obtained after performing model selection over its hyperparameter space $\Theta$, using grid-search against a validation set. The validation set is a label-stratified subset (a $20\%$ split) of the current labeled set; the rest is used for training.
  2. 2.
    The model is also calibrated [^1] A notable exception is in our use of the RoBERTa model, which already is well calibrated [9].. This is critical since query strategies $\mathcal{Q}$ often use the predicted class probabilities from $M_{t}$. We use Platt scaling [26, 24].
The process is initialized by selecting a seed set of size $s$ from the unlabeled data pool, using a strategy $\mathcal{A}$ (line 2). We use random selection for this step.
We also note that a “model” here might mean a combination of a text representation, e.g., word vectors, and a classifier, e.g., Random Forest; further detailed in Section 4.
It might seem unusual that we explicitly mention model selection and calibration, which should be part of the standard practice for building probabilistic classifiers. Unfortunately, in our experience multiple implementations/studies miss one or both of these steps.
4 Comparison Methodology
In our experiments, we vary classifiers, text representations (we often jointly refer to them as a prediction pipeline), batch size, seed size and, of course, query strategies. These combinations are listed below:
  1. 1.
    Prediction pipeline: There are two categories of pipelines we use:
    1. (a)
      Separate representation and classifier: The representations used are USE [3], MPNet [31] and word vectors (we use the models provided by the spaCy library [15]). While USE and MPNet directly convert sentences into vectors, in the case of word vectors, we compute the average of the vectors of all words in a text to obtain its representation. For classification, we useRandom Forests (RF) [2] and Support Vector Machines [6] with a linear kernel - we’ll term the latter as “LinearSVC”.
      We use off-the-shelf representations and they are not fine-tuned on our data. Only the classifiers are trained on our data.
    2. (b)
      End-to-end classifier: This does not require a separate representation model. We use RoBERTa [17] (a variant of BERT). This is fine-tuned on the labeled data at each AL iteration.
    Hyperparameter search spaces are detailed in Section 0.A.4.2 of the Appendix. As noted in Section 3, model selection and calibration are performed during training of a prediction pipeline. The only exception is RoBERTa, which has been shown to be well-calibrated out of the box [9].
    The first category gives us $2\times 3=6$ combinations. Counting RoBERTa, we have $\boldsymbol{7}$ prediction pipelines in our study.
  2. 2.
    Query Strategy: we list these below, with the year of publication mentioned, to show our focus on contemporary techniques:
    1. (a)
      Random: the batch is selected uniformly at random. This forms our baseline.
    2. (b)
      Margin [^2] Also referred to as Smallest Margin or Breaking Ties. [27] (2001): this selects instances with the smallest differences between the confidence of the most likely and the second-most likely predicted (by the current classifier [^3] Note that in reference to Algorithm 1, at iteration $t$, the query strategy $\mathcal{Q}$ uses model $M_{t-1}$.) classes. Despite being a relatively old technique, it continues to be competitive [28].
    3. (c)
      Contrastive Active Learning (CAL) [21] (2021): chooses instances whose predicted class-probability distribution is the most different (based on KL divergence) from those of their $k$-nearest neighbors. This is similar to another work [23], where such conflicts are detected using the explanation space produced by XAI techniques.
    4. (d)
      Discriminative Active Learning (DAL) [13, 11] (2019): a binary classifier (a feedforward neural network) is constructed to discriminate between labeled and unlabeled data, and then selects unlabeled instances with the greatest predicted probability of being unlabeled. This picks examples that are most different from the labeled instances in this classifier’s representation space. While the original work [13] only considers image datasets, a separate study shows its efficacy on text [11].
    5. (e)
      Representative Errors for Active Learning (REAL) [5] (2023): identifies clusters in the unlabeled pool and assigns the majority predicted label as a “pseudo-label” to all points in it. Instances are then sampled whose predictions differ from the pseudo-label. The extent of disagreement and cluster size are factored into the sampling step
    We use a total of $\boldsymbol{5}$ query strategies.
  3. 3.
    Datasets: we use $\boldsymbol{5}$ standard datasets: agnews, sst-2, imdb, pubmed and dbpedia-5 (a $5$-label version of the standard dbpedia dataset that we created). These are detailed in Table 2 in the Appendix.
  4. 4.
    Batch and Seed sizes: We use batch and seed size combinations of $(200,200)$ and $(500,500)$. This is a total of $\boldsymbol{2}$ combinations.
  5. 5.
    Trials: For statistical significance, we run $\boldsymbol{3}$ trials for each combination of the above settings.
Other settings: The accuracy metric we use is the F1 (macro) score, since it prevents performance on a dominant class from overwhelming results. The size of the unlabeled pool is $20000$ at the start of each experiment. If the original dataset has more than than $20000$ instances, we extract a label-stratified sample, to retain the original class distribution. The size of the test set is $5000$ also a label-stratified sample from the corresponding test set of the original dataset.
We run an experiment till the size of the labeled set has grown to $5000$ instances. This which implies $T=(5000-200)/200=24$ iterations for the batch/seed size setting of $(200,200)$, and similarly $9$ iterations for the $(500,500)$ setting.
Total # experiments [^4] Our code is available here: https://anonymous.4open.science/r/On_the_Fragility_of_Active_Learners/README.md.: Accounting for the above settings, we have $7\times 5\times 5\times 2=350$ unique configurations. Accounting for the trials, we have a total of $350\times 3=\boldsymbol{1050}$ experiments. Reiterating something we had mentioned earlier: we follow the due process of performing model selection and calibration (excepting RoBERTa) in each AL iteration of each of these experiments.
5 Reproducibility of Benchmarked Techniques
Before we analyze our results, we would like to stress that our intention is not to suggest that the techniques we evaluate, e.g., REAL, CAL, DAL, don’t work. In the specific settings discussed in their respective papers, they most likely perform as reported. In the interest of fairness, we have conducted limited independent tests that confirm this - see Section 0.A.3 in the Appendix.
6 Results
We present our observations in this section. We begin with visualizing a cross-section of our experiments: Figure 1 shows results for the dataset agnews and batch/seed size of $(200,200)$.
Figure 1: F1 macro scores on the test set at each iteration, for the dataset agnews and batch size of $200$. The $x$-axes show size of the labeled data, the $y$-axes show the F1-macro scores on the test data. The corresponding plot for batch size = $500$ is shown as Figure 8 in the Appendix.
We introduce some notation to aid in the interpretation of our results. We think of the F1-macro score to be provided by this function:
$f(g,\mathcal{Q},\mathcal{D},b,s,N)$ (1)
Here,
  • $g$ denotes a prediction pipeline.
  • $\mathcal{Q}\in\{cal,dal,real,margin,random\}$ is a query strategy.
  • $\mathcal{D}\in\{imdb,agnews,dbpedia-5,pubmed,sst-2\}$ is a query strategy.
  • $b\in\{200,500\}$ is the batch size.
  • $s\in\{200,500\}$ is the seed size.
  • $N$ is the size of the labeled data. Here $N\leq 5000$.
[…]
We now try to answer some interesting questions based on our results. In most cases, we will average over various factors, especially datasets, since most AL techniques assume their query strategy may be used without any prior knowledge.
6.1 In what data regimes do query strategies most differ?
We immediately note in Figure 1 that learning curves vary quite a bit based on the prediction pipeline.
What broad trend(s) might we expect? F1-macro scores from different query strategies (for a given pipeline and dataset) should converge as we see more data due to at least two reasons: (a) the concept space would be eventually covered, (b) at later iterations, there is less of the unlabeled pool to choose from. Indeed, Figure 2 confirms this. We first compute variances in F1-macro scores for different pipeline/dataset combinations [^5] This step comes first since the accuracies obtained by a LinearSVC would be very different from those by RoBERTa. In this case, the inner quantity is conditioned over the dataset first, because we wanted to validate a specific property. In subsequent analyses, we will average over datasets, so as to mimic real-life use of AL, where no prior knowledge is assumed., and then plot their averages. In other words, a point $(x,y)$ for the batch/seed size $(200,200)$ curve denotes:
$y=\mathbb{E}_{g,\mathcal{D}}[Var[f|g=g^{\prime},\mathcal{D}=d^{\prime},N=x,b=200,s=200]]$ (2)
where $g^{\prime}$ and $d^{\prime}$ are a specific pipeline and dataset respectively.
Figure 2: Expectation over variance of F1-macro given a pipeline and dataset, plotted size of labeled data. See Section 6.1 for details. Note that the batch size doesn’t strongly affect trends.
6.2 Measuring effectiveness
A key reason to use AL is its effectiveness over random sampling. A common way to assess this is to use the Wilcoxon signed-rank test, e.g., [11], by comparing accuracy scores from these strategies, at various labeled data sizes. Unfortunately, this entirely ignores the magnitude of the difference in scores. This is a limitation of rank based metrics in general [^6] Another example is Mean Rank over AUC values of the learning curves.. See Figure 3, which shows learning curves for the random strategy, and query strategies $QS1$ and $QS2$. The latter two strategies have identical p-values of $9.77\times 10^{-4}$ for the Wilcoxon test (compared to random).
Figure 3: Hypothetical query strategies $QS1$ and $QS2$ shown against random. Note that p-value from the Wilcoxon test is identical for both. Discussed in Section 6.2.
In practice, one would clearly prefer $QS2$ over $QS1$, since it reaches higher accuracies earlier. In fact, considering the possibly non-zero cost of setting up an AL-based tagging system, a query strategy needs to be significantly better for it to be useful.
We recommend assessing the magnitude (relative to the random strategy) using Root Mean Square Error (RMSE) or the Area Under the (Learning) Curve (AUC).
RMSE vs AUC: the choice between them comes down to these factors: (a) RMSE doesn’t provide a direction of change (while AUC does), which may be addressed by pairing it with the Wilcoxon test, (b) AUC monotonically increases from left-to-right on the $x$-axis (unlike RMSE which is a mean), and so it must be carefully interpreted. Direct comparisons of AUC are most convenient over identical spans of the $x$-axis.
We apply these metrics to our results to obtain Figure 4, where we show both average p-values from the Wilcoxon test (on the left) and average RMSE scores (right subplot). Mathematically, for a given value on the $x$-axis, we compare the sequences:
$(\mathbb{E}_{g,\mathcal{D}}[f|\mathcal{Q}\in S_{qs},N=n,b=200,s=200])_{n\leq x}\text{ and}$
$(\mathbb{E}_{g,\mathcal{D}}[f|\mathcal{Q}=random,N=n,b=200,s=200])_{n\leq x}$ (3)
where $S_{qs}=\{cal,dal,real,margin\}$. Note that the $y$ values are cumulative, i.e., all sizes $n\leq x$ are considered. We prefer the RMSE here to normalize for the $x$-axis span.
For the Wilcoxon test, we have set the minimal sample size at $3$ points, which is why we see different starting points on the $x$-axes in Figure 4. This test is also performed to assess if the alternative, i.e., a non-random strategy, produces higher values than the random strategy.
Figure 4: The left plot shows averaged p-values from a Wilcoxon test, while the right plot shows averaged RMSE values. A consistently good query strategy should have a low p-value and high RMSE; none of the evaluate strategies fall in that category. Discussed in Section 6.2.
An ideal query strategy should have a low p-value and high RMSE. Unfortunately, no query strategy fulfills this criteria: while cal and dal have high RMSE values, they have poor p-values (the dal curve eclipses cal in the plot). And while real and margin have p-values that improve with labeled data, they exhibit small RMSE scores. Some things of note:
  1. 1.
    The fact the p-value scores of real and margin improve with size, is something we will encounter again (Figure 5).
  2. 2.
    The RMSE plot reinforces trends of the variance plot in Figure 2: we see higher RMSEs in the low data regime.
  3. 3.
    The batch/seed settings preserve the general trend.
6.3 Gain over random and margin strategies
We now ask if we actually do better than random sampling. We define the relative improvement $\delta$ between two scores as:
$\delta(treatment,control)=100\times\frac{(treatment-control)}{control}$ (4)
For a labeled dataset of size $(n_{1},n_{2}]$, batch/seed size $(b^{\prime},s^{\prime})$ and
$q^{\prime}\in\{cal,dal,real,margin\}$, we compute:
$\mathbb{E}_{g,\mathcal{D}}[\delta((f|\mathcal{Q}=q^{\prime},b=b^{\prime},s=s^{\prime},n_{1}<N\leq n_{2}]),\\ (f|\mathcal{Q}=random,b=b^{\prime},s=s^{\prime},n_{1}<N\leq n_{2}))]$ (5)
This is shown in Figure 5, where the labeled data size is broken up into multiple bins to distinguish patterns at different size regimes. We also separate the batch sizes (see legend). Given the negative scores, it seems that in general, random is either competitive or better!
At relatively higher data sizes, margin and real start to surpass random: which seems to suggest that (a) at low size regimes, where we expect the greatest impact from AL, their performance is actually poor, and (b) after gathering some data, these strategies are able to better exploit structure in the data (and margin seems surprisingly competitive in this regime). Also, recall that we had seen this behaviour of margin and real in the RMSE subplot of Figure 4.
Since behaviour across the two batch/seed size settings don’t seem to appreciably differ (we note this in Figures 2, 4 and 5), we will average over them in the next few plots.
Figure 5: Average relative improvement in F1-macro score over random, for different query strategies. Corresponds to Equation 5.
We can repeat the above exercise, to assess the effect of pipelines now, i.e., for a given data size bin $(n_{1},n_{2}]$ and prediction pipeline $g^{\prime}$, we now compute:
$\mathbb{E}_{\mathcal{Q}\in S_{qs},\mathcal{D},b,s}[\delta((f|g=g^{\prime},n_{1}<N\leq n_{2}]),\\ (f|g=g^{\prime},\mathcal{Q}=random,n_{1}<N\leq n_{2}))]$ (6)
where $S_{qs}=\{cal,dal,real,margin\}$. This is plotted in Figure 6. Here we note that, using RoBERTa gives us a better average performance over random at any size regime.
Figure 6: Average relative improvement in F1-macro score over random, for different pipelines. This corresponds to Equation 6.
We might want to know if for the promising case of RoBERTa relatively new techniques like dal, real and cal outperform margin. This is easy to plot with minor changes to the quantity in Equation 6:
$\mathbb{E}_{\mathcal{Q}\in S_{qs},\mathcal{D},b,s}[\delta((f|g=RoBERTa,n_{1}<N\leq n_{2}]),\\ (f|g=RoBERTa,\mathcal{Q}=margin,n_{1}<N\leq n_{2}))]$ (7)
where $S_{qs}=\{cal,dal,real\}$. The corresponding plot is shown in Figure 7. We see that aside from one data regime, $(1250-2500]$ margin remains quite competitive.
Figure 7: Average relative improvement in F1-macro score over margin for RoBERTa. This corresponds to Equation 7.
6.4 The best combinations
Finally, we ask what are the best combinations of pipelines and query strategies, at different size regimes? We measure this by the AUC per regime. Given a pipeline $g^{\prime}$, strategy $q^{\prime}$, and data sizes $n_{i}$ within a bin, the AUC [^7] Using AUC is reasonable since the $x$-axis is broken up into bins of roughly equal size (see discussion in Section 6.2), and we are not looking for comparison to the random strategy. is computed for the following sequence:
$(\mathbb{E}_{\mathcal{D},b,s}[f|g=g^{\prime},\mathcal{Q}=q^{\prime},N=n_{i}])_{n_{i}}$ (8)
Table 1 shows the top-$10$ combinations of pipelines and query strategies for different regimes. Not surprisingly, wrt pipelines, RoBERTa shows the strongest performance. Somewhat surprisingly, the next best pipeline is LinearSVC-MPNet (shortened to LinSVC_MPNet) and not a RF-based pipeline. Our guess as to why this is the case is that RF’s hyperparamṡearch space (detailed in Section 0.A.4.2) was not comprehensive enough, while MPNet transformed the data to be nearly linearly separable, which enabled the high-bias LinearSVC to quickly find a good optima. Sadly, no single query strategy dominates the QS column, and thus we are offered no insight as to what strategy pairs best with a pipeline. Somewhat noteworthy is the fact that the random strategy is quite competitive for the LinearSVC-MPNet pipeline, but not for RoBERTa.
Table 1: The top-$10$ combinations of pipelines and query strategies are shown (in two columns) for different labeled data size regimes. We show the AUC score for a combinations in size regime. See Section 6.4 for details.
train size pipeline QS AUC pipeline QS AUC
(0-1250] 1. RoBERTa dal 875.46 6. LinSVC_MPNet random 807.04
(0-1250] 2. RoBERTa real 873.64 7. LinSVC_MPNet real 794.69
(0-1250] 3. RoBERTa margin 873.36 8. LinSVC_MPNet margin 793.08
(0-1250] 4. RoBERTa cal 871.92 9. LinSVC_MPNet cal 790.42
(0-1250] 5. RoBERTa random 862.99 10. LinSVC_MPNet dal 789.60
(1250-2500] 1. RoBERTa cal 986.22 6. LinSVC_MPNet random 913.76
(1250-2500] 2. RoBERTa dal 984.40 7. LinSVC_MPNet real 907.98
(1250-2500] 3. RoBERTa real 981.07 8. LinSVC_MPNet margin 907.39
(1250-2500] 4. RoBERTa margin 976.07 9. LinSVC_MPNet dal 900.13
(1250-2500] 5. RoBERTa random 974.04 10. LinSVC_MPNet cal 899.12
(2500-3750] 1. RoBERTa margin 903.60 6. LinSVC_MPNet random 837.91
(2500-3750] 2. RoBERTa dal 900.93 7. LinSVC_MPNet margin 836.57
(2500-3750] 3. RoBERTa real 900.83 8. LinSVC_MPNet real 833.95
(2500-3750] 4. RoBERTa cal 897.16 9. LinSVC_MPNet dal 828.73
(2500-3750] 5. RoBERTa random 891.58 10. LinSVC_MPNet cal 827.24
(3750-5000] 1. RoBERTa real 1087.92 6. LinSVC_MPNet margin 1016.94
(3750-5000] 2. RoBERTa margin 1087.37 7. LinSVC_MPNet random 1010.88
(3750-5000] 3. RoBERTa dal 1085.06 8. LinSVC_MPNet real 1008.67
(3750-5000] 4. RoBERTa cal 1076.79 9. LinSVC_MPNet dal 1004.44
(3750-5000] 5. RoBERTa random 1075.20 10. LinSVC_MPNet cal 999.97
7 Summary and Conclusion
We first summarize our observations:
  1. 1.
    We noted in the dataset cross-section (Figure 1) of our experiments that the behavior of query strategies vary widely wrt pipelines. However, we observe the larger expected trend of reduction in the variance of accuracies produced by different strategies, with increase in the size of labeled data (Figure 2 in Section 2).
  2. 2.
    Wrt assessing the value of a query strategy, we noted that the Wilcoxon test only provides a partial picture by ignoring magnitudes (Figure 3 in Section 6.2). We might use the RMSE score to measure differences wrt the random strategy. Applied to our results, it shows that no query strategy exhibits the favorable properties of having both a low p-value and high RMSE score (Figure 4 in Section 6.2).
  3. 3.
    Our batch/seed size settings of $(200,200)$ and $(500,500)$ only seem to influence slight change in behaviors. We noted this in Figure 2, Figure 4 and Figure 5.
  4. 4.
    Most query strategies exhibit no significant gain over the random strategy (Figure 5), although at relatively higher data regimes we noted real and margin surpass random.
  5. 5.
    Among pipelines, RoBERTa exhibits the strongest performance (Figure 6) relative to random, but not wrt margin (Figure 7).
  6. 6.
    When we list the top-$10$ best performing combinations per data size regime (Table 1), we note that RoBERTa and LinearSVC-MPNet consistently do well, but unfortunately, their best performances are paired with different query strategies. The only (broad) takeaways are that while LinearSVC-MPNet works well with the random strategy, RoBERTa performs well with non-random strategies, but its not clear which specific one.
After extensive evaluation of different AL algorithms, we are forced to conclude that AL doesn’t reliably work across datasets, prediction pipelines and text representations. This study may be considered “limited” relative to real-world variances, e.g., many more choices of classifiers, datasets, which leads us to suspect that the true picture is probably more dismal. Of course, we see some associations in isolated pockets, such as RoBERTa consistently performs well, that are promising, but not enough (with what query strategy?) to derive a useful prescription. It’s known that query strategies require a “warm-start” labeled set, and this is probably why we observe the deferred performance of real and margin in Figure 5, but again, this does not translate to practical advice unless we can precisely identify a settings-specific warm-start size. It’d seem the only way to know if an AL technique would work in a setting is to find results reported for similar settings, and even this advice leaves open the question of inferring similarity between datasets. And if one can’t find such results, the only other way to know is by trying out multiple techniques - which of course is a chicken-and-egg problem when data-labeling is expensive!
How do we encourage the field of AL to be more utility-driven? We believe there are at least two ways: (a) embrace a broader discourse where the success of a technique needs to be tied to fundamental properties of datasets (e.g., topological features [4]) and predictors (e.g., VC dimension [32]), that are identifiable in an unsupervised manner in novel settings, and (b) on the empirical front, reported results need to account for an array of settings much more diverse than is customary today.
Appendix 0.A Appendix
0.A.1 Datasets
As an addendum to the methodology discussion in Section 4, we provide details of the datasets used in Table 2.
Table 2: Datasets used.
Dataset
# classes
Description
sst-2
2
Single sentences extracted from movie reviews with their sentiment label [30].
imdb
2
Movie reviews with corresponding sentiment label [19].
agnews
4
News articles with their topic category [36].
pubmed
4
Sentences in medical articles’ abstracts which are labeled with their role on the abstract [8].
dbpedia-5
5
A subset of dbpedia [36] which contains Wikipedia articles accompanied by a topic label. The original dataset’s instances are evenly distributed across $14$ classes. To form dbpedia-5, we use only the first $5$ classes: Company, EducationalInstitution, Artist, Athlete, OfficeHolder. This was done to reduce the training time of one-vs-all classifiers, e.g., LinearSVC.
0.A.2 Learning curves for agnews
This is the plot for the batch/seed size setting of $(500,500)$, for the dataset agnews. This is a companion to the plot shown in Figure 1 of Section 6.
Figure 8: F1 macro scores on the test set at each iteration, for the dataset agnews and batch size of $500$. The $x$-axes show size of the labeled data, the $y$-axes show the F1-macro scores on the test data.
0.A.3 Reproducibility Experiments
In all cases, we have attempted to replicate the original settings, e.g. same train/development/test data split, model type, seed/batch sizes, number of AL iterations as shown in Table 3. For CAL, REAL, we report the F1-macro scores on agnews, in which classes are evenly distributed, instead of the accuracy provided in the original papers. For DAL, we use the dataset cola [^8] https://nyu-mll.github.io/CoLA/ and utilise the Hugging Face library to finetune BERT (while the original work employs TensorFlow [^9] https://www.tensorflow.org/, but we use equivalent settings). Figure 9 shows a comparison between our results and the reported ones in these papers. Despite some minor differences in the setups, we observe that these AL methods work as described in their respective papers in these settings.
One significant difference between these settings compared to our methodology is the use of a predetermined labeled development set for all BERT/RoBERTa model finetuning. This set is relatively larger than the AL batch or seed size and is not part the labeled data available at each AL iteration. This is impractical in scenarios where AL is typically used: labeling is expensive. Moreover, in some cases, there is no model selection performed, which we remedy in our experiments (Section 3).
Figure 9: Comparison between published results and ours with the same settings for CAL, REAL, DAL.
Table 3: Settings for reproducibility experiments.
AL
Dataset
AL loop
Classifier & text representation
QS parameters
Metric
CAL
agnews
b=2280
s=1140
T = 7
BERT (bert-base-cased)
[CLS] at the last hidden layer
learning rate = 2e-5
train batch size = 16
# epochs = 3
sequence length = 128
warmup ratio = 0.1
# evaluations per epoch = 5
# neighbors=10
F1-macro
DAL
cola
b=50
s=100
T = 5
BERT (bert-base-uncased)
[CLS] at the pooled layer
learning rate = 5e-5
train batch size = 50
# epochs = 5
sequence length = 50
warmup ratio = 0
# evaluations per epoch = 1
Accuracy
REAL
agnews
b=150
s=100
T = 8
RoBERTa (roberta-base)
[CLS] at the last hidden layer
learning rate = 2e-5
train batch size = 8
# epochs = 4
sequence length = 96
warmup ratio = 0.1
# evaluations per epoch = 4
# clusters=25
F1-macro
0.A.4 Hyperparameters
0.A.4.1 Query Strategy (QS) hyperparameters
For each QS’s hypeparameters, we use the values recommended by the authors in corresponding papers. This means setting number of nearest neighbors in CAL to 10, number of clusters in DAL to 25, and keeping the same discriminative model in REAL.
0.A.4.2 Hyperparameters search for prediction pipelines
Table 4 shows the search space for hyperparameters we use for each classifier.
Table 4: Hyperparameters for each classifier in the prediction pipelines.
Classifier
Hyperparameters
RoBERTa
roberta-base
[CLS] at the last hidden layer
learning rate = {3e-5, 5e-5}
train batch size = 16
# epochs = {5, 10}
sequence length = 128
warmup ratio = 0.1
# evaluations per epoch = 5
LinearSVC
C = {0.001, 0.01, 0.1, 1, 10, 100, 1000}
class weight = balanced
RF
min samples leaf = {1, 5, 9}
# estimators = {5, 10, 20, 30, 40, 50}
max depth = {5, 10, 15, 20, 25, 30}
class weight = balanced
max features = sqrt